Julia's soybean field is 3m longer than it is wide. To increase production, she plans to increase both the length and width by 2m. If the new field is 46m^2 larger than the old field, then what are the dimension of the old field?

Answer by TimothyLamb(4379)

---
x = length
y = width
---
x = y + 3
---
area of old field:
a = xy
---
area of new field:
A = (x + 2)(y + 2)
---
A = a + 46
(x + 2)(y + 2) = xy + 46
xy + 2x + 2y + 4 = xy + 46
2x + 2y = 42
---
system of linear equations:
x = y + 3
2x + 2y = 42
---
put the system of linear equations into standard form
---
x - y = 3
2x + 2y = 42
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
answer:
x = length = 12 m
y = width = 9 m
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php